OpenStudy (anonymous):

the sum of six and four times d is the same as d minus nine. translate each sentence into an equation.

OpenStudy (anonymous):

6+4d=d-9

OpenStudy (anonymous):

thank you

OpenStudy (lilg132):

6 + 4d = d - 9 4d - d = -9 + 6 3d = -3

OpenStudy (lilg132):

d = -1